[Ncep.nhc.nco_contacts] HSU Shift Log -- Technical Edition -- Sat Oct 28, 2017

HSU Shift Log nhc.tsbadmin at noaa.gov
Sat Oct 28 21:05:10 UTC 2017


HSU Shift Log for Sat Oct 28 2017 17:05:03 GMT-0400 (EDT)

Shift: H

Forecaster: SR Stewart

Technical Issues Summary:

ATCF issues: None.

N-AWIPS issues: Radar data from any US radar site or radar composite has  
been unavailable during the entire shift. Using Internet and AWIPS.

AWIPS-2/NCP issues: None.

Other Techincal issues: None.

Miscellaneous issues: None.
-------------- next part --------------
An HTML attachment was scrubbed...
URL: https://www.lstsrv.ncep.noaa.gov/pipermail/ncep.nhc.nco_contacts/attachments/20171028/39b8af86/attachment.html 


More information about the Ncep.nhc.nco_contacts mailing list